The atoms of an LiCl molecule are separated by a distance r = 0.200 nm.

(a) Calculate the reduced mass of an LiCl molecule.
kg

(b) Calculate the moment of inertia of an LiCl molecule.
kg · m2

(c) Calculate the wavelength of radiation emitted when an LiCl molecule undergoes a transition from the J = 2 state to the J = 1 state.
cm
Answer #1

(a)

\mu ={\frac {m_{1}m_{2}}{m_{1}+m_{2}}}

where u is reduced mass

u = 6*35/(6+35)

u = 5.121

(b) MOment of inertia I = m*M*r^2/(m+M)

I = 6*35*(0.2*10^(-9))^2/(41)

= 2.04*10^-20 kgm^2